OpenStudy (anonymous):

Suppose oil spills from a ruptured tanker and spreads in a circular pattern. If the radius of the oil spill increases at a constant rate of 2.5 m/s. How fast is the area of the spill increasing when the radius is 14 m?

OpenStudy (bahrom7893):

dr/dt = 2.5 m/s r = 14 m dA/dt - ?

OpenStudy (bahrom7893):

dA/dt = d(pi*r^2)/dt = 2*pi*r*(dr/dt)

OpenStudy (bahrom7893):

Plug in everything you have: dA/dt = 2*pi*2.5*14 = 5*14pi = 70pi